# Quantum Quasinormal Mode Theory for Dissipative Nano-Optics and Magnetodielectric Cavity Quantum Electrodynamics

### v1

@@ IV Effective Lindblad master equation
V Conclusion
@@
In conclusion, we have introduced a formalism for the quantization of QNMs in the presence of 3D, spatially inhomogeneous, dissipative (with possible gain), linear magnetodielectric media, enabling a unified and rigorous approach for simulating photonic, plasmonic and magnonic cQED phenomena. Similar to a number of recent works Franke et al. (2019); Medina et al. (2021); Feist, Fernández-Domínguez, and García-Vidal (2020), our approach is rooted in the phenomenological Green’s function quantization scheme of mQED, but relies on the use of PMLs to achieve a set of modes that are rigorously regularized, orthogonalized and complete, rendering them suitable for the representation of the dyadic Green’s function and electromagnetic fields both inside and outside the resonator volume. The introduction of PMLs transforms the radiative losses into non-radiative material dissipation, and via a suitable transformation that reflects all the losses of the resonator, we are able to define creation and annihilation operators that allow the construction of modal Fock states for the joint excitations of field-dressed matter. Our quantized description of the QNMs thereby provides the foundations for a quantum dynamic treatment of magnetodielectric cQED systems, utilizing discrete modes together with rigorously derived quantum Langevin and master equations. By directly addressing the intricacies of modal loss in a fully quantum theory of magnetodielectric cQED, our approach enables a proper assessment of the performance metrics for nanophotonic device protocols in the Quantum 2.0 era, extending from on-demand sources of single photons, plasmons and magnons, dynamic entanglement generation and quantum state transduction, to light-matter strong coupling in hybrid configurations. Moreover, our quantum QNM theory could prove especially valuable in revealing the opportunities that dissipation may offer in manipulating and tailoring the dynamics of quantum systems (contrary to its traditionally perceived adversarial role), such as in the development of quantum technologies based on quantum nanoplasmonic coherent perfect absorption Lai et al. (2024), non-hermitian topological magnonics Yu et al. (2024) and dissipative photonic time crystals Lyubarov et al. (2022); Liu et al. (2023); Sustaeta-Osuna, García-Vidal, and Huidobro (2025). Note that although the modification of mQED in the case of time-dependent material properties is not trivial, important first steps have very recently been made with the proposition of an mQED Lagrangian for a time-dependent Drude model Horsley and Baker (2025).

### v2

@@ IV Effective Lindblad master equation
V Numerical Example: 1D Half-open Cavity
Having presented the general formalism in the preceding sections, we now demonstrate the predictive capabilities of our approach in a simple, 1D example. As schematized in Fig. 3(a), we consider a half-open cavity comprising a homogeneous, diamond sample of relative permittivity $\varepsilon=5.7$ and length $L=200\,\textrm{nm}$, with a perfect electrical conductor boundary placed at $x=0$ (corresponding to a Dirichlet boundary condition imposed at this point). We assume that the diamond hosts a single QE at some location $x_{\mathrm{QE}}$ along the $x$ axis, featuring a pair of levels that are coupled by the electric dipole interaction with the cavity field. Such a QE could be realized in the form of a color center in the diamond lattice, and a variety of such optically active defects have been explored for quantum applications Aharonovich and Neu (2014). For the purpose of our numerical tests, we shall here consider a QE having the transition energy of a nitrogen-vacancy center, namely $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$ ($637\,\mathrm{nm}$), and a vertically oriented transition dipole moment ${\bm{d}}$ with controllable magnitude $|{\bm{d}}|$.
A 1D formulation of mQED can be derived for systems that are invariant in two spatial dimensions (say, the $y$ and $z$ directions), and can be used to study the interaction of QEs with linearly polarized radiation propagating in the $x$ direction Gruner and Welsch (1996). In this case, we make the replacements $\hat{\bm{E}}(\bm{r})\rightarrow\hat{E}(x)=\hat{E}_{y}(x)$, ${\bm{d}}\rightarrow d=d_{y}$ and $\hat{\bm{f}}(\bm{r},\omega)\rightarrow\hat{f}(x,\omega)=\hat{f}_{y}(x,\omega)$. The $y$ component of the noise current density is then given by
$\hat{j}_{\mathrm{F}}(x,\omega)=\hat{j}_{\mathrm{F},y}(x,\omega)$ $=\omega\sqrt{\frac{\hbar\varepsilon_{0}}{\pi A}\tilde{\varepsilon}(x,\omega)}\hat{f}_{\mathrm{e}}(x,\omega)$ (83)
$+\partial_{x}\left(\sqrt{-\frac{\hbar}{\pi\mu_{0}A}\tilde{\kappa}(x,\omega)}\hat{f}_{\mathrm{m}}(x,\omega)\right)\ ,$
in which $A$ is the normalization area in the $yz$ plane, and we adopt 1D material parameters $\tilde{\zeta}(x,\omega)=\tilde{\zeta}_{yy}(x,\omega)$ appropriate to a single PML in the $x$ direction, given by
$\tilde{\zeta}(x,\omega)=\zeta(x,\omega)\left(1+\frac{\mathrm{i}}{\omega+\mathrm{i}\eta}\sigma_{x}(x)\right)\ ,$ (84)
as can be obtained from Eqs. (9) and (10) with $\sigma_{y,z}=0$. In the 1D case, the PMLs are purely lossy, as waves with normal incidence experience only their loss components. In our example, we choose PMLs with a step-like absorption profile, $\sigma_{x}(x)=2.5\,\mathrm{eV}\cdot\theta(x-L)$ (i.e., they are directly adjacent to the resonator), a thickness of $20L$ and $\eta=0.1\,\mathrm{eV}$. It is worth noting that this 1D problem admits an analytical solution for the QNMs and Green’s function directly from the boundary and continuity conditions, thus obviating the need for numerical solution of the 1D Helmholtz equation. As such, spurious numerical reflections that would otherwise arise from a spatial discretization do not occur, and a simple, step-like absorption profile, rather than a gradually increasing one, suffices.
The formulae for the $P$-matrices all have the same form as in the 3D case, but with the replacements mentioned above and ×$\nabla\times\rightarrow\partial_{x}$:
$P_{nn^{\prime}}=\frac{\hbar\mu_{0}}{\pi A}\int_{0}^{\mathrm{\infty}}\mathrm{d}\omega\,\omega^{2}A_{n}(\omega)A_{n^{\prime}}^{*}(\omega)$ (85)
$\int\mathrm{d}s\,\frac{\omega^{2}}{c^{2}}\tilde{\varepsilon}_{\mathrm{i}}(s,\omega)\tilde{E}_{n}(s)\tilde{E}^{*}_{n^{\prime}}(s)-\tilde{\kappa}_{\mathrm{i}}(s,\omega)\partial_{s}\tilde{E}_{n}(s)\partial_{s}\tilde{E}^{*}_{n^{\prime}}(s)\ ,$
where we have combined the electric and magnetic modal operators. The corresponding master equation then has the same form as in Eq. (81) with $\tilde{\chi}_{nn^{\prime}}^{\mathrm{PML+G},-}=0$ and $\tilde{P}_{nn^{\prime}}^{\mathrm{L}}=P_{nn^{\prime}}$. In the weak coupling or bad cavity limit, the electromagnetic modes can be traced out to obtain a reduced master equation for the QE only, as was done in Refs. Franke et al. (2019); Cirac (1992), with the QE decay rate
$\Gamma_{\mathrm{QE}}=\frac{1}{\hbar^{2}}\Re\left(g_{n}P^{+\frac{1}{2}}_{nn^{\prime}}\frac{1}{\mathrm{i}(\omega_{\mathrm{QE}}-\tilde{\omega}_{n^{\prime}}^{*})}P^{-\frac{1}{2}}_{n^{\prime}n^{\prime\prime}}g_{n^{\prime\prime}}^{*}\right)\ .$ (86)
Normalizing this result with respect to the free-space decay rate in 1D, $\Gamma_{0}={\omega_{\mathrm{QE}}|d|^{2}/}{2\hbar n_{\mathrm{b}}\varepsilon_{0}cA}$, yields the Purcell factor.
For our calculations using Eq. (86), we retain the three lowest-order QNMs of the resonator which lie in spectral proximity to the QE transition $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$, with complex frequencies $\tilde{\omega}_{1}=(0.649-0.184\mathrm{i})\,\mathrm{eV}$, $\tilde{\omega}_{2}=(1.95-0.184\mathrm{i})\,\mathrm{eV}$ (i.e., resonant with the QE transition) and $\tilde{\omega}_{3}=(3.25-0.184\mathrm{i})\,\mathrm{eV}$, as shown in the upper panel of Fig. 3(b). These modes have very low quality factors of $Q_{1}=1.76$, $Q_{2}=5.28$ and $Q_{3}=8.80$. Furthermore, we assume that the QE is located at $x_{\mathrm{QE}}=66\,\mathrm{nm}$, where an electric-field antinode of the second-order QNM occurs [see Fig. 3(a)]. Under these conditions, whilst all three QNMs of the resonator offer polarization matching, only the second-order QNM allows optimal spectral and spatial matching with the QE transition. Fig. 3(b) presents our calculated Purcell factor, which bears the expected structure comprising three Lorentzian-like peaks at the spectral locations of the QNMs. In the weak coupling regime considered here, the decay rate can be derived semi-classically using Fermi’s golden rule and the Green’s function of the electric field Novotny and Hecht (2006), yielding (in 1D) $\Gamma^{\textrm{SC}}_{\mathrm{QE}}=\omega_{\mathrm{QE}}^{2}|d|^{2}\mathrm{Im}(G(x_{\mathrm{QE}},x_{\mathrm{QE}},\omega_{\mathrm{QE}}))/\hbar\varepsilon_{0}c^{2}A$. The predictions obtained using the fully quantum and semi-classical models are compared in the lower panel of Fig. 3(b), showing excellent agreement. Note that we have considered only a finite frequency interval $[\omega_{\mathrm{min}},\omega_{\mathrm{max}}]$ for the numerical integration in Eq. (85), with $\omega_{\mathrm{min}}=0.25\,\mathrm{eV}$ and $\omega_{\mathrm{max}}=4.0\,\mathrm{eV}$.
In the next step, we compare the electric field and QE dynamics in the strong coupling regime. In Ref. Dung, Knöll, and Welsch (2000); Scheel and Buhmann (2008), the spontaneous decay of QEs was studied within the framework of mQED. If the QE is prepared in the excited state |$\ket{\mathrm{e}}$ at time $t=0$, and the electromagnetic field in the vacuum state |$\ket{0}$, then the time-dependent state vector of the system can be written as |$\ket{\psi(t)}=C_{\mathrm{e}}(t)\mathrm{e}^{-\mathrm{i}\omega_{\mathrm{QE}}t}\ket{\mathrm{e}}\ket{0}+\int\mathrm{d}x\int\mathrm{d}\omega\,C_{\mathrm{g},\lambda}(x,\omega,t)\mathrm{e}^{-\mathrm{i}\omega t}\ket{\mathrm{g}}\hat{f}_{\lambda}^{\dagger}(x,\omega)\ket{0}$. From the Heisenberg equations of mQED, an integral equation for the time-dependent coefficient $C_{\mathrm{e}}(t)$ can be derived:
$C_{\mathrm{e}}(t)=1+\int_{0}^{\mathrm{t}}\mathrm{d}\tau\,K(t-\tau)C_{\mathrm{e}}(\tau)\ ,$ (87)
with the kernel
$K(t-\tau)=\frac{|d|^{2}}{\hbar\varepsilon_{0}\pi A}\int_{\omega_{\mathrm{min}}}^{\omega_{\mathrm{max}}}\mathrm{d}\omega\,\frac{\omega^{2}}{c^{2}}\frac{\Im(G(x_{\mathrm{QE}},x_{\mathrm{QE}},\omega))}{\mathrm{i}(\omega-\omega_{\mathrm{QE}})}$
$\cdot\left(\mathrm{e}^{-\mathrm{i}(\omega-\omega_{\mathrm{QE}})(t-\tau)}-1\right)\ .$ (88)
This equation is derived from the exact mQED formulation and involves no approximations beyond the rotating-wave one for the interaction with the QE. We have solved the integral equation numerically, considering once more a QE tuned to the second-order QNM (i.e., $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$) and setting the dimensionless dipole moment to be $|d|/\sqrt{\hbar\varepsilon_{0}cA}=0.3$, which although unrealistically high, serves to demonstrate the validity of our theory in the strong coupling case. The result for the occupation probability of the excited state, $|C_{\mathrm{e}}(t)|^{2}$, can then be compared to the expectation value $\langle\sigma^{+}(t)\sigma^{-}(t)\rangle$, obtained from the quantum QNM master equation that we solve using the QuTiP library Lambert et al. (2024). The numerical data are plotted in Fig. 4(a), evidencing excellent agreement between the exact theory and the quantum QNM approach in capturing the temporal Rabi oscillations characteristic of this coupling regime. We have further plotted the expectation values of the QNM number operators $\langle\hat{a}^{\dagger}_{n}(t)\hat{a}_{n}(t)\rangle$; as expected, the symmetrized, second-order mode is dominant and undergoes coherent energy exchange with the QE, as reflected in the Rabi oscillations. Moreover, the amplitudes decrease in an approximately linear fashion on the logarithmic plot, with a gradient corresponding to the (equal) imaginary parts of the QNM frequencies $\Im(\tilde{\omega}_{n})=-0.184\,\mathrm{eV}$. After the first Rabi oscillation, the other two symmetrized modes show qualitatively similar oscillations as the second-order one, which might be explained by the contribution of the second-order QNM in the other symmetrized modes. Prior to this (i.e., during the first $\sim 10\,\mathrm{fs}$), their dynamics is more complicated and shows additional structure caused by the excitation of the first- and third-order QNMs.
From the exact mQED formulation, one can further derive an equation for the expectation value of the electric field intensity emitted via the spontaneous decay of the QE Dung, Knöll, and Welsch (2000):
$\langle\hat{E}^{(-)}(x,t)\hat{E}^{(+)}(x,t)\rangle=\bigg{|}\frac{d}{\pi\varepsilon_{0}c^{2}\sqrt{A}}\int_{0}^{t}\mathrm{d}\tau\,C_{\mathrm{e}}(\tau)$
$\cdot\int_{\omega_{\mathrm{min}}}^{\omega_{\mathrm{max}}}\mathrm{d}\omega\,\omega^{2}\Im(G(x,x_{\mathrm{QE}},\omega_{\mathrm{QE}}))\mathrm{e}^{-\mathrm{i}(\omega-\omega_{\mathrm{QE}})(t-\tau)}\bigg{|}^{2}\ .$ (89)
In Fig. 4(b), we compare the exact result for the normalized output intensity at $x=L$ with that obtained from the QNM master equation; once again, very good agreement can be seen. We observe two peaks of the intensity attributable to the Rabi oscillations of the second-order mode around $\sim 5\,\mathrm{fs}$ and $\sim 17\,\mathrm{fs}$. However, the first peak has additional substructure. Its maximum is split into a pair of maxima and the second derivative is changing sign along the drop of the peak. For comparison, we plot the intensity calculated from the quantum QNM model upon neglecting the first- and third-order modes (i.e., $|\tilde{E}^{\mathrm{sym}}_{2}(x)|^{2}\langle\hat{a}^{\dagger}_{2}(t)\hat{a}_{2}(t)\rangle$). From this, we can appreciate that the additional substructure originates from a multi-mode effect, as the intensity in this case only displays the two peaks from the Rabi oscillations.
Collectively, our results show that the derived master equation correctly reproduces the dynamics of the QE and electric field, seeding confidence in its application to more complex, quantum nanophotonic systems.
It is noteworthy that in this 1D example, we can place the PMLs directly at the open end of the resonator. Positioning the PMLs away from the open end would degrade the accuracy of the results, and increasingly so the farther away they are shifted, given the divergence and incompleteness of the QNMs outside the resonator. In such circumstances, a correct expansion of the Green’s function in this region would necessitate taking the PML modes into account. For 3D systems, where there is a transition between the near-field and the divergent far-field regions, optimizing the separation between the resonator and PMLs should aid in reducing the required number of PML modes, a matter to be investigated in greater detail in future studies.
Figure 3: Numerical exploration of a 1D, half-open, diamond cavity in air, with an electric-dipole QE embedded therein. (a) Schematic of the cavity-QE system with characteristic parameters indicated. The dashed line shows the electric field intensity profile of the second-order QNM, to which the QE is resonantly coupled. (b) Complex frequencies $\tilde{\omega}_{n}$ of the three, lowest-order QNMs (upper panel) and the frequency-dependent Purcell factor of a QE with $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$ ($637\,\mathrm{nm}$), calculated in accordance with the quantum QNM and semi-classical theories (lower panel).
Figure 4: Spontaneous decay of a QE with $\omega_{\mathrm{QE}}=1.95\,\mathrm{eV}$ and $|d|/\sqrt{\hbar\varepsilon_{0}cA}=0.3$ in the 1D, half-open cavity of Fig. 3(a). (a) QE and symmetrized mode dynamics obtained by solving the QNM master equation, compared to the QE dynamics obtained from the integral equation Eq. (87) in mQED. (b) Normalized output electric field intensity at $x=L$ obtained from the QNM master equation, compared to the exact mQED result from Eq. (89). Also shown is the electric field intensity calculated from the quantum QNM model retaining only the second-order mode $\hat{a}_{2}$.
VI Conclusion
@@
In conclusion, we have introduced a formalism for the quantization of QNMs in the presence of 3D, spatially inhomogeneous, dissipative (with possible gain), linear magnetodielectric media, enabling a unified and rigorous approach for simulating photonic, plasmonic and magnonic cQED phenomena. Similar to a number of recent works Franke et al. (2019); Medina et al. (2021); Feist, Fernández-Domínguez, and García-Vidal (2020), our approach is rooted in the phenomenological Green’s function quantization scheme of mQED, but relies on the use of PMLs to achieve a set of modes that are rigorously regularized, orthogonalized and complete, rendering them suitable for the representation of the dyadic Green’s function and electromagnetic fields both inside and outside the resonator volume. The introduction of PMLs transforms the radiative losses into non-radiative material dissipation, and via a suitable transformation that reflects all the losses of the resonator, we are able to define creation and annihilation operators that allow the construction of modal Fock states for the joint excitations of field-dressed matter. Our quantized description of the QNMs thereby provides the foundations for a quantum dynamic treatment of magnetodielectric cQED systems, utilizing discrete modes together with rigorously derived quantum Langevin and master equations. By directly addressing the intricacies of modal loss in a fully quantum theory of magnetodielectric cQED, our approach enables a proper assessment of the performance metrics for nanophotonic device protocols in the Quantum 2.0 era, extending from on-demand sources of single photons, plasmons and magnons, dynamic entanglement generation and quantum state transduction, to light-matter strong coupling in hybrid configurations. Moreover, our quantum QNM theory could prove especially valuable in revealing the opportunities that dissipation may offer in manipulating and tailoring the dynamics of quantum systems (contrary to its traditionally perceived adversarial role), such as in the development of quantum technologies based on quantum nanoplasmonic coherent perfect absorption Lai et al. (2024), non-hermitian topological magnonics Yu et al. (2024) and dissipative photonic time crystals Lyubarov et al. (2022); Liu et al. (2023); Sustaeta-Osuna, García-Vidal, and Huidobro (2025). Note that although the modification of mQED in the case of time-dependent material properties is not trivial, important first steps have very recently been made with the proposition of an mQED Lagrangian for a time-dependent Drude model Horsley and Baker (2025), following the quantization of the electromagnetic field in lossless systems Pendry and Horsley (2024).